# Two-step estimation of latent trait models

### v2

@@ 3.2 Asymptotic properties and variance estimation
In practical data analysis, we could also choose to omit $\mathbf{V}_{1}$ from the variance. This would mean, in effect, that once step 1 was done we would be treating the measurement model obtained from it, and thus the definition of the latent variables $\boldsymbol{\eta}_{i}$, as known and fixed rather than as an estimable characteristic. Estimated structural models obtained from step 2 would then be models for this definition of latent variables. This is close in spirit to the approach that is also almost always adopted in naive three-step estimation, as discussed further in Section 3.4.2 below. This may be desirable or unproblematic in many applications. When it is not, however, omitting the contribution from $\mathbf{V}_{1}$ may result in serious underestimation of the variance of $\tilde{\boldsymbol{\theta}}_{2}$.
[…]
3.3 Implementation of the estimation
Software that can do one-step estimation of a model can also be used for two-step estimation of it. We have used Mplus 6.12 software (Muthén and Muthén, 2010) for the simulation studies and application example of this paper. We have supplemented it with functions in R (R Core Team, 2022) to manage the estimation process. This automatically sequences the two steps of estimation, passing the estimated parameter values from the first step to the code for the second and the estimates back to R. We have used the MplusAutomation package in R (Hallquist and Wiley, 2018) to control Mplus from R, and the brew package (Horner, 2011) to automatically edit the input files. Examples of the estimation code are included in supplementary materials for the paper.
Because the log likelihood functions can have multiple local maxima, it is desirable to use multiple starting values for the estimation algorithm. This is a standard feature in software such as Mplus. High-quality software is also needed to implement and control the numerical integration that is needed in all parts of the estimation. These numerical challenges apply to both two-step and one-step estimation of latent trait models, although the starting point issue may be somewhat smaller for the simpler two-step approach.
[…]
@@ 3.4.1 One-step estimation
The main alternative to two-estimation is one-step estimation. It means obtaining maximum likelihood (ML) estimates $\hat{\boldsymbol{\theta}}=(\hat{\boldsymbol{\theta}}_{1},\hat{\boldsymbol{\theta}}_{2})$ of the measurement parameters and structural parameters together, by maximizing the log-likelihood (9) with respect to $\boldsymbol{\theta}$. This can be done using the same software (in our applications Mplus) that can also be used for two-step estimation.
@@ 3.4.2 Naive three-step estimation
Three-step estimation is a stepwise approach. Its first step is the same as in the two-step method, i.e. obtaining estimates $\tilde{\boldsymbol{\theta}}_{1}$ of the measurement parameters from simplest possible models. In the second step, these are used to calculate predicted values $\tilde{\eta}_{ik}$ for the latent variables, for each $i=1,\dots,n$; $k=1,\dots,K$. In the third step, the structural model is estimated with $\tilde{\boldsymbol{\eta}}_{i}=(\tilde{\eta}_{i1},\dots,\tilde{\eta}_{iK})^{\prime}$ substituted for $\boldsymbol{\eta}_{i}$ and treated as observed variables. For the recursive model (2)–(4) this means simply fitting the $M$ regression models for $\tilde{\boldsymbol{\eta}}_{i(1)},\dots,\tilde{\boldsymbol{\eta}}_{i(M)}$ separately. Estimated variances of the parameter estimates are also obtained as in standard regression, in effect treating $\tilde{\eta}_{ik}$ like any observed variables. We refer to this version of the approach as “naive” three-step estimation.
The predictions (or “latent trait scores”) $\tilde{\eta}_{ik}$ could be obtained in different ways. The most common, and the ones we use here, are the conditional expected values (empirical Bayes predictions) $\tilde{\eta}_{ik}=\mbox{E}(\eta_{k}|\mathbf{Y}_{i(k)};\tilde{\boldsymbol{\psi}}_{(k)})$, where $\mathbf{Y}_{i(k)}$ denotes those items which were included in step-1 estimation of the measurement model for $\eta_{ik}$ and $\tilde{\boldsymbol{\psi}}_{(k)}$ the estimated parameters of this model. These $\tilde{\eta}_{ik}$ require numerical integration, but they are normally provided by estimation software.
This approach is “naive” because in most cases it produces biased estimates of the parameters of the structural model. This is because $\tilde{\eta}_{ik}$ are erroneously measured versions of $\eta_{ik}$. Trying to correct for this bias leads to “adjusted” versions of three-step estimation, which we discuss in the next section. Before that, however, it is worth noting that whether we would want to do that depends on what we think the goal of three-step estimation is. Measurement error bias is a problem if we want to treat the step-3 models as estimates of the regression models (2)–(4) for the latent variables $\boldsymbol{\eta}_{i}$. But it is not a problem if we view them instead as estimates for the same models for the scores $\tilde{\boldsymbol{\eta}}_{i}$ themselves. In other words, we can choose to treat $\tilde{\boldsymbol{\eta}}_{i}$ as variables of interest rather than as poor measures of $\boldsymbol{\eta}_{i}$. In this view, steps 1 and 2 of the analysis become data reduction steps which are used only to develop a calculation that reduces $\mathbf{Y}_{i}$ to these $\tilde{\boldsymbol{\eta}}_{i}$. This also an entirely coherent approach to data analysis, if it is substantively interesting and justified. Here, however, we focus on the situation where models for the latent $\boldsymbol{\eta}_{i}$ are the goal, because it is also what motivates two-step and one-step estimation.
@@ 3.4.3 Adjusted three-step estimation
[…]
Note that here we deviate slightly from previous notation, to make one point clearer below. This is that the latent variables are denoted not as $\boldsymbol{\eta}_{i}$, but as $\boldsymbol{\eta}=(\eta_{.1},\dots,\eta_{.K})^{\prime}$ to emphasise the fact that in (14) they are integration variables and not fixed values for units $i$. Adjusted three-step estimation could be based on the right-hand side or left-hand side of (14), and this has been done for other families of latent variable models. For latent trait models, however, this approach is less promising, because it is either redundant or unavailable as a general method.
Consider first the right-hand side of (14). This is still a latent variable model, but with $\tilde{\boldsymbol{\eta}}_{i}$ rather than $\mathbf{Y}_{i}$ as the observed measurement indicators. It is thus like step 2 of two-step estimation, but applied to $\tilde{\boldsymbol{\eta}}_{i}$. To implement it, we would first need to derive the form of the distribution $p(\tilde{\boldsymbol{\eta}}_{i}|\boldsymbol{\eta})$. This is fairly easy for latent class models, where $\mathbf{Y}_{i}$, $\boldsymbol{\eta}$ and $\tilde{\boldsymbol{\eta}}_{i}$ are all categorical, and estimation based on this approach has been developed there by Vermunt (2010) and Bakk et al. (2013). It could also be straightforward for normal structural equation models (SEMs), but it has not been examined in full in that literature (perhaps because other three-step methods discussed below are then more natural and straightforward). Savalei (2018) considered it for SEMs in the special case where $\tilde{\eta}_{ik}$ are equally-weighted sum scores of the measurement items.
To examine the situation for latent trait models, suppose for simplicity that all the items $Y_{ikj}$ are binary and that the measurement model is (7), i.e. that the different traits $\eta_{\cdot k}$ are measured by the distinct sets of items $\mathbf{Y}_{ik}=(Y_{ik1},\dots,Y_{ikp_{k}})^{\prime}$. We can then write (14) as
[…]
@@
There is, however, a different way of obtaining these three-step estimates. Consider again binary items, and suppose that all estimated measurement models (8) for different items $Y_{ikj}$ for $\eta_{.k}$ are different from each other. The weights $a_{kj}$ in the approximation $\tilde{\eta}_{ik}\approx\sum_{j}a_{kj}Y_{ikj}$ are then also all different, so that every possible value of $\mathbf{Y}_{ik}$ gives a distinct value of $\tilde{\eta}_{ik}=\tilde{\eta}_{ik}(\mathbf{Y}_{ik})$ almost surely. When this is the case, $p(\tilde{\eta}_{ik}|\eta_{.k})=p(\mathbf{Y}_{ik}|\eta_{.k})$ and $p(\tilde{\boldsymbol{\eta}}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})$ in (15) is equal to the $p(\mathbf{Y}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})$ that would be used for step 2 of two-step estimation. It would seem that the same argument holds also for the exact values of $\tilde{\boldsymbol{\eta}}_{i}$, and when any items $Y_{ikj}$ are polytomous. In other words, if all the items have different estimated measurement parameters, there is almost surely a one-to-one relationship between distinct values of $\mathbf{Y}_{i}$ and $\tilde{\boldsymbol{\eta}}_{i}$. Adjusted three-step estimation based on the right-hand side of (14), if correctly implemented, would then give the same estimates of $\boldsymbol{\theta}_{2}$ as two-step estimation which uses the same step-1 estimates of $\boldsymbol{\theta}_{1}$ — but would deliver them with more effort for no gain. If different items had identical measurement models (most likely because they were constrained to be so), this kind of three-step estimation would require separate implementation, but this would again be effectively pointless because it would in the end reflect essentially the same information that two-step estimation uses more easily. We do not consider such three-step estimates further in this paper.
@@
For latent trait models this would require a sufficiently good approximation of the integral which would also yield a conveniently estimable model. That is not generally available. In some cases, however, a version of it can be derived. Suppose that the interesting part of the structural model is $\eta_{2i}=\beta_{20}+\beta_{21}\eta_{1i}+\psi_{2i}$, with no covariates $\mathbf{X}_{i}$. Suppose further that we use for $\eta_{1i}$ the Empirical Bayes score $\tilde{\eta}_{1i}=\mbox{E}(\eta_{.1}|\mathbf{Y}_{i1})$, and for $\eta_{2i}$ a unbiased score $\breve{\eta}_{2i}$ for which $\mbox{E}(\breve{\eta}_{2i}|\eta_{2i})=\eta_{2i}$. Then we have $\breve{\eta}_{2i}=\beta_{20}+\beta_{21}\tilde{\eta}_{1i}+\psi_{2i}^{*}$, so that $(\beta_{20},\beta_{21})$ can be estimated from a linear regression model for $\tilde{\eta}_{2i}$ given $\tilde{\eta}_{1i}$ (which has a heteroscedastic residual $\psi_{2i}^{*}$, but this can be allowed for in the estimation). For this model and these trait scores, naive and adjusted three-step estimation are thus the same. This situation was described for SEMs by Skrondal and Laake (2001), using regression factor scores as $\tilde{\eta}_{1i}$ and Bartlett factor scores as $\breve{\eta}_{2i}$. Lu and Thomas (2008) showed that the same construction would work also for latent trait models, but they also noted that for these models there are no unbiased scores to be used as $\breve{\eta}_{2i}$ (the best that can be obtained is consistency as the number of items increases).
@@
This approach would not generalise well to more complex models. For example, suppose that the structural model includes a further latent variable $\eta_{3i}$ and the model $\eta_{3i}=\beta_{30}+\beta_{31}\eta_{1i}+\beta_{32}\eta_{2i}+\psi_{3i}$. The scores that should be assigned for $(\eta_{1i},\eta_{2i})$ to estimate this model would be $\mbox{E}[(\eta_{1i},\eta_{2i})|\mathbf{Y}_{1i},\mathbf{Y}_{2i}]$, i.e. predictions given both of their indicators (Skrondal and Laake 2001, Lu and Thomas 2008); these should also be conditional on any covariates $\mathbf{X}_{i}$ in the structural model. To get these, we would need to fit the step-1 model for $\eta_{1i}$ and $\eta_{2i}$ jointly rather than separately. Furthermore, the appropriate trait scores to assign for $\eta_{1i}$ and $\eta_{2i}$ would be different for the model for $\eta_{3i}$ than for the model for $\eta_{2i}$, and different again for any further parts of the structural model. This could quickly get very impracticable for larger models. We do not consider this version of adjusted three-step estimation further in this paper, except for some examples with only a single explanatory latent trait in the simulation studies below.

### v3

@@ 3.2 Asymptotic properties and variance estimation
Di Mari and Kuha (2025) proposed an alternative, simulation-based way of estimating $\mathbf{V}$ which avoids the need to evaluate (12). It is based on the decomposition $\mbox{var}(\tilde{\boldsymbol{\theta}}_{2})=\mbox{E}_{\tilde{\boldsymbol{\theta}}_{1}}[\mbox{var}(\tilde{\boldsymbol{\theta}}_{2}|\tilde{\boldsymbol{\theta}}_{1})]+\mbox{var}_{\tilde{\boldsymbol{\theta}}_{1}}[\mbox{E}(\tilde{\boldsymbol{\theta}}_{2}|\tilde{\boldsymbol{\theta}}_{1})]$ where the outer expectation and variance are over the sampling distribution of the step-1 estimator $\tilde{\boldsymbol{\theta}}_{1}$. The first term of this is estimated by the same $\hat{\mathbf{V}}_{2}$ as above. The second term can be estimated by drawing multiple values of $\boldsymbol{\theta}_{1}$ from the sampling distribution of $\tilde{\boldsymbol{\theta}}_{1}$, carrying out step-2 estimation given each of them in turn, and calculating the variance matrix of the resulting sample of estimates of $\boldsymbol{\theta}_{2}$. This is asymptotically equivalent with variance estimation from (12). Essentially the same idea is employed by Levy (2023) and Levy and McNeish (2025) in their Bayesian approach, which uses draws of the parameters from MCMC simulations.
@@
A rather different approach would be to omit $\mathbf{V}_{1}$ from the variance altogether. This would mean, in effect, that once step 1 was done we would treat the measurement model obtained from it, and thus the definition of the latent variables $\boldsymbol{\eta}_{i}$, as known and fixed rather than as an estimable characteristic. Estimated structural models obtained from step 2 would then be models for this definition of the latent variables. This is close in spirit to the approach that is also almost always adopted in naive three-step estimation, as discussed further in Section 3.4.2 below. This may be desirable or unproblematic for practical data analysis in many applications. When it is not, however, omitting the contribution from $\mathbf{V}_{1}$ may result in serious underestimation of the variance of $\tilde{\boldsymbol{\theta}}_{2}$.
Consider, finally, the potential deviation from the conditions of PML estimation that was mentioned above. It is a distributional inconsistency which can arise when the latent variables $\boldsymbol{\eta}_{i}$ are continuous. We assume throughout this paper that the conditional distributions for $\boldsymbol{\eta}_{i}$ are normal, as shown in (3)–(4). This also implies that $p(\mathbf{\eta}_{ik}|\mathbf{X}_{i})$ is univariate normal for all $k$. The marginal distribution of any single latent trait is then $p(\eta_{ik})=\int p(\eta_{ik}|\mathbf{X}_{i})p(\mathbf{X}_{i})\,d\mathbf{X}_{i}$ where $p(\mathbf{X}_{i})$ is the joint sample distribution of $\mathbf{X}_{i}$. This $p(\eta_{ik})$ is not normal unless $p(\mathbf{X}_{i})$ is also normal. The marginal normal distribution that we assume in (10) for $\eta_{ik}$ (and any multivariate normal distributions for multiple traits which may be used in step 1) is thus misspecified to some extent, if the structural model includes non-normal covariates $\mathbf{X}_{i}$. Literature on latent trait models with misspecified distributions of the trait (see e.g. Manapat and Edwards, 2022, and references therein) indicates that estimates of the measurement parameters can then be biased, with a bias that depends on the true distribution of $\eta_{ik}$ and is largest when it is very skewed. In two-step estimation this would be a concern if such bias in $\tilde{\boldsymbol{\theta}}_{1}$ from step 1 also translated into bias in $\tilde{\boldsymbol{\theta}}_{2}$ from step 2. We examine this situation in the simulations in Section 4. There we do not observe any meaningful bias of this kind in $\tilde{\boldsymbol{\theta}}_{2}$, even when the true trait distribution in step 1 is very non-normal. Taking the latent traits to be normally distributed also in step 1 thus seems empirically justified. We note also that the same approximation is routinely made also in one-step estimation when we consider different choices of $\mathbf{X}_{i}$ and take $\eta_{ik}$ be normally distributed given any of them (which cannot be exactly true for all of them).
3.3 Model selection
We have described how two-step estimation will be done for any given model specification. In practice, however, we first need to decide which specification(s) to use for substantive analysis. This too is best done in two steps, selecting first the form of the measurement model and then the structural model (this is in fact an older definition of “two-step analysis” in latent variable modelling; see Anderson and Gerbing 1988). The basic tools for doing this are the standard methods of likelihood-based model comparison, such as likelihood ratio tests and the penalised model selection criteria AIC and BIC.
Measurement models would again be first examined with the structural model omitted as far as possible, and typically further split up into smaller pieces. Here, where the substantive interest is on structural models, it should be rare for us to want to examine the measurement properties of all the indicators together, in a fully exploratory manner (if that was still thought necessary, it is likely that the understanding of the measurement items was not yet good enough to allow meaningful substantive analysis of models between the latent constructs). What is then left to examine, if anything, are smaller choices within the measurement models, such as possible conditional associations between items, cross-loadings between indicators of closely related constucts, or aberrant behaviour of individual items.
Once the form of the measurement model is selected and its parameters fixed at their estimated values, different structural models can be compared as a routine part of step 2 of two-step estimation. Here we would, in particular, examine which (latent and observed) explanatory variables need or need not be included in different parts of the structural model. These steps are essentially similar as they would be for standard regression models with all variables observed.
Elsewhere in latent variable modelling, especially in structural equation modelling, it is also common to carry out “global” assessments of a whole model at once, rather than its individual parameters separately. This can use overall significance tests (“overidentification tests”) or goodness-of-fit indices (many of which exist, especially for SEMs). In a two-step context this would be much less natural, because it would be a retrograde step: global model assessment again blends measurement and structural models, which the two-step approach takes pains to separate. We thus do not recommend such global methods here. In principle, however, they could still be implemented (see Rosseel and Loh 2024 for more on this). For any indices this would be done simply by calculating them as normal using the two-step parameter estimates, whereas for goodness-of-fit tests their sampling distributions might need to be modified.
@@ 3.4.2 Naive three-step estimation
Three-step estimation is another stepwise approach. Its first step is the same as in the two-step method, i.e. obtaining estimates $\tilde{\boldsymbol{\theta}}_{1}$ of the measurement parameters from simplest possible models. In the second step, these are used to calculate predicted values $\tilde{\eta}_{ik}$ for the latent variables, for $i=1,\dots,n$; $k=1,\dots,K$. In the third step, the structural model is estimated with $\tilde{\boldsymbol{\eta}}_{i}=(\tilde{\eta}_{i1},\dots,\tilde{\eta}_{iK})^{\prime}$ substituted for $\boldsymbol{\eta}_{i}$ and treated as observed variables. For the recursive model (2)–(4) this means simply fitting the $M$ regression models for $\tilde{\boldsymbol{\eta}}_{i(1)},\dots,\tilde{\boldsymbol{\eta}}_{i(M)}$ separately. Estimated variances of the parameter estimates are also obtained as in standard regression, in effect treating $\tilde{\eta}_{ik}$ like any observed variables. We refer to this as “naive” three-step estimation.
The predictions (or “latent trait scores”) $\tilde{\eta}_{ik}$ could be obtained in different ways. The ones we use here are the conditional expected values (empirical Bayes predictions) $\tilde{\eta}_{ik}=\mbox{E}(\eta_{k}|\mathbf{Y}_{i(k)};\tilde{\boldsymbol{\psi}}_{(k)})$, where $\mathbf{Y}_{i(k)}$ denotes those items which were included in step-1 estimation of the measurement model for $\eta_{ik}$ and $\tilde{\boldsymbol{\psi}}_{(k)}$ the estimated parameters of this model. These $\tilde{\eta}_{ik}$ require numerical integration, but they are normally provided by estimation software.
This approach is “naive” because $\tilde{\eta}_{ik}$ are erroneously measured versions of $\eta_{ik}$. As a result, in most cases it produces biased estimates of the parameters of the structural model. Trying to correct for this bias leads to “adjusted” versions of three-step estimation, which we discuss in the next section. Before that, however, it is worth noting that whether we would want to do that depends on what we think the goal of three-step estimation is. Measurement error bias is a problem if we want to treat the step-3 models as estimates of the regression models (2)–(4) for the latent variables $\boldsymbol{\eta}_{i}$. But it is not an issue if we view them instead as estimates for the same models for the scores $\tilde{\boldsymbol{\eta}}_{i}$ themselves. In other words, we can choose to treat $\tilde{\boldsymbol{\eta}}_{i}$ as variables of interest rather than as poor measures of $\boldsymbol{\eta}_{i}$. In this view, steps 1 and 2 of the analysis become data reduction steps which are used only to develop a calculation that reduces $\mathbf{Y}_{i}$ to these $\tilde{\boldsymbol{\eta}}_{i}$. This also an entirely coherent approach to data analysis, if it is substantively interesting and justified. Here, however, we focus on the situation where models for the latent $\boldsymbol{\eta}_{i}$ are the goal, because it is also what motivates two-step and one-step estimation.
@@ 3.4.3 Adjusted three-step estimation
$p(\tilde{\boldsymbol{\eta}}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})=\int\,p(\tilde{\boldsymbol{\eta}}_{i}|\boldsymbol{\eta})\,p(\boldsymbol{\eta}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})\,d\boldsymbol{\eta}.$ (14)
Note that here we deviate slightly from previous notation, to make one point clearer below. This is that the latent variables are denoted not as $\boldsymbol{\eta}_{i}$, but as $\boldsymbol{\eta}=(\eta_{.1},\dots,\eta_{.K})^{\prime}$ to emphasise the fact that in (14) they are integration variables and not fixed values for units $i$. Adjusted three-step estimation could be based on the right-hand side or left-hand side of (14), and this has been done for other families of latent variable models. For latent trait models, however, it is less promising, because it is either redundant or unavailable as a general method.
Consider first the right-hand side of (14). This is still a latent variable model, but with $\tilde{\boldsymbol{\eta}}_{i}$ rather than $\mathbf{Y}_{i}$ as the observed measurement indicators. It is thus like step 2 of two-step estimation, but applied to $\tilde{\boldsymbol{\eta}}_{i}$. To implement it, we would first need to derive the form of the distribution $p(\tilde{\boldsymbol{\eta}}_{i}|\boldsymbol{\eta})$. This is fairly easy for latent class models, where $\mathbf{Y}_{i}$, $\boldsymbol{\eta}$ and $\tilde{\boldsymbol{\eta}}_{i}$ are all categorical, and estimation based on this approach has been developed there by Vermunt (2010) and Bakk et al. (2013). It could also be straightforward for normal structural equation models (SEMs), but it has not been examined in full in that literature (other three-step methods discussed below are more natural and straightforward for them). Savalei (2018) considered it for SEMs in the special case where $\tilde{\eta}_{ik}$ are equally-weighted sum scores of the measurement items.
To examine this possibility for latent trait models, suppose for simplicity that all the items $Y_{ikj}$ are binary and that the measurement model is (7), i.e. that the different traits $\eta_{\cdot k}$ are measured by the distinct sets of items $\mathbf{Y}_{ik}=(Y_{ik1},\dots,Y_{ikp_{k}})^{\prime}$. We can then write (14) as
$p(\tilde{\boldsymbol{\eta}}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})=\int\,\left[\prod_{k=1}^{K}p(\tilde{\eta}_{ik}|\eta_{\cdot k})\right]\,p(\boldsymbol{\eta}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})\,d\boldsymbol{\eta}.$ (15)
@@
There is, however, a different way of obtaining these three-step estimates. Consider again binary items, and suppose that all estimated measurement models (8) for different items $Y_{ikj}$ for $\eta_{.k}$ are different from each other. The weights $a_{kj}$ in the approximation $\tilde{\eta}_{ik}\approx\sum_{j}a_{kj}Y_{ikj}$ are then also all different, so that every possible value of $\mathbf{Y}_{ik}$ gives a distinct value of $\tilde{\eta}_{ik}=\tilde{\eta}_{ik}(\mathbf{Y}_{ik})$ almost surely. When this is the case, $p(\tilde{\eta}_{ik}|\eta_{.k})=p(\mathbf{Y}_{ik}|\eta_{.k})$ and $p(\tilde{\boldsymbol{\eta}}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})$ in (15) is equal to the $p(\mathbf{Y}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})$ that would be used for step 2 of two-step estimation. It would seem that the same argument holds also for the exact values of $\tilde{\boldsymbol{\eta}}_{i}$, and when any items $Y_{ikj}$ are polytomous. In other words, if all the items have different estimated measurement parameters, there is almost surely a one-to-one relationship between distinct values of $\mathbf{Y}_{i}$ and $\tilde{\boldsymbol{\eta}}_{i}$. Adjusted three-step estimation based on the right-hand side of (14), if correctly implemented, would then give the same estimates of $\boldsymbol{\theta}_{2}$ as two-step estimation which uses the same step-1 estimates of $\boldsymbol{\theta}_{1}$ — but would deliver them with more effort for no gain. If different items were constrained to have identical measurement models, this kind of three-step estimation would require separate implementation, but this would again be effectively pointless because it would in the end reflect essentially the same information that two-step estimation uses more easily. We do not consider such three-step estimates further in this paper.
@@
For latent trait models this would require a sufficiently good approximation of the integral which would also yield a conveniently estimable model. That is not generally available. In some cases, however, a version of it can be derived. Suppose that the interesting part of the structural model is $\eta_{2i}=\beta_{20}+\beta_{21}\eta_{1i}+\psi_{2i}$, with no covariates $\mathbf{X}_{i}$. Suppose further that we use for $\eta_{1i}$ the Empirical Bayes score $\tilde{\eta}_{1i}=\mbox{E}(\eta_{.1}|\mathbf{Y}_{i1})$, and for $\eta_{2i}$ an unbiased score $\breve{\eta}_{2i}$ for which $\mbox{E}(\breve{\eta}_{2i}|\eta_{2i})=\eta_{2i}$. Then we have $\breve{\eta}_{2i}=\beta_{20}+\beta_{21}\tilde{\eta}_{1i}+\psi_{2i}^{*}$, so that $(\beta_{20},\beta_{21})$ can be estimated from a linear regression model for $\tilde{\eta}_{2i}$ given $\tilde{\eta}_{1i}$ (which has a heteroscedastic residual $\psi_{2i}^{*}$, but this can be allowed for in the estimation). For this model and these trait scores, naive and adjusted three-step estimation are thus the same. This situation was described for SEMs by Skrondal and Laake (2001), using regression factor scores as $\tilde{\eta}_{1i}$ and Bartlett factor scores as $\breve{\eta}_{2i}$. Lu and Thomas (2008) showed that the same construction would work also for latent trait models, but they also noted that for these models there are no unbiased scores to be used as $\breve{\eta}_{2i}$ (the best that can be obtained is consistency as the number of items increases).
@@
This approach would not generalise well to more complex models. For example, suppose that the structural model includes a further latent variable $\eta_{3i}$ and the model $\eta_{3i}=\beta_{30}+\beta_{31}\eta_{1i}+\beta_{32}\eta_{2i}+\psi_{3i}$. The scores that should be assigned for $(\eta_{1i},\eta_{2i})$ to estimate this model would be $\mbox{E}[(\eta_{1i},\eta_{2i})|\mathbf{Y}_{1i},\mathbf{Y}_{2i}]$, i.e. predictions given both of their indicators (Skrondal and Laake 2001, Lu and Thomas 2008); these should also be conditional on any covariates $\mathbf{X}_{i}$ in the structural model. To get these, we would need to fit the step-1 model for $\eta_{1i}$ and $\eta_{2i}$ jointly rather than separately. Furthermore, the appropriate trait scores to assign for $\eta_{1i}$ and $\eta_{2i}$ would be different for the model for $\eta_{3i}$ than for the model for $\eta_{2i}$, and different again for any further parts of the structural model. This would quickly get very impracticable for larger models. We do not consider this version of adjusted three-step estimation further in this paper, except for some examples with only a single explanatory latent trait in the simulation studies below.
3.5 Implementation of the estimation
Software that can do one-step estimation of a model can also be do two-step estimation of it. We have used Mplus 6.12 software (Muthén and Muthén, 2010) for the simulations and application example of this paper. We have supplemented it with functions in R (R Core Team, 2022) to manage the process. This automatically sequences the two steps of estimation, passing the estimated parameter values from the first step to the code for the second and the estimates back to R. We have used the MplusAutomation package in R (Hallquist and Wiley, 2018) to control Mplus from R, and the brew package (Horner, 2011) to automatically edit the input files. Examples of the estimation code are included in supplementary materials for the paper.
The integrals in the log likelihoods (9) and (10) have no closed form, so numerical integration is needed to evaluate them. This is the most demanding computational element of likelihood-based estimation of these models, and it is important that high-quality software is used to implement and control it. The computational demands are substantially smaller for two-step than for one-step estimation, both because step 1 typically involves only one-dimensional integrals and because step 2 tends to require fewer iterations (and hence fewer evaluations of the integrals). For example, most of the computing time in our simulations in Section 4 was spent on one-step estimation, and there were similarly large differences in our substantive application in Section 5, as discussed further there. It is likely that this is the most important practical difference between these two methods in large models, where one-step estimation quickly becomes more and more demanding. For both approaches, the computational task gets harder with more latent variables. In our examples we have considered models with up to three latent variables without problems, but much larger models would be correspondingly more challenging.
Because the log likelihoods can have multiple local maxima, it is desirable to use multiple starting values for the estimation. This is a standard feature in software such as Mplus.
3.6 Other methods: Non-likelihood-based estimation
The focus of this paper is on a likelihood-based framework, where the one- and two-step methods are full-information ML and pseudo-ML estimation respectively. Other methods of estimation are also available for latent trait models. For example, in the SEM literature it is common to consider limited-information methods of the kind proposed by Muthén (1984) (see also Muthén and Satorra 1995) such as Diagonally Weighted Least Squares (DWLS) estimation. They in effect convert models with categorical indicators into SEMs for continuous indicators, under some further assumptions. This is done by presenting the observed categorical indicators $\mathbf{Y}$ as coarsened versions of continuous latent variables $\mathbf{Y}^{*}$ that follow a multivariate normal distribution given $\mathbf{X}$, using the observed data to estimate the means and variance matrix of $\mathbf{Y}^{*}$ given $\mathbf{X}$, and then treating these as data for estimating the model of interest with some form of weighted least squares estimation. This produces estimates that are often nearly as efficient as one-step FIML estimates. These methods have some limitations of applicability, arising from the assumption of normality for $\mathbf{Y}^{*}$. For example, they cannot be used for models with unordered categorical indicators, or with the logit models for binary or ordinal ones that we use (they imply probit models instead). Their handling of missing data is also somewhat inflexible. Their main advantage is computation, which can be very substantially easier than for likelihood-based estimation.
With reference to our focus, standard use of limited-information estimation like this is still one-step estimation, in that both measurement and structural parameters are estimated together in its last stage. As such, comparisons between full- and limited-information estimates are outside the scope of this paper. However, it is worth noting that the choice between these methods of estimation is separate from the choice of one or two steps in the estimation. The two-step idea can be combined with any methods of estimation which satisfy very general conditions (see Gourieroux and Monfort, 1995, Sections 24.2.4 and 24.2.2), and it will still give estimates that are consistent and asymptotically normal. Applied to DWLS, say, this would mean carrying out its last stage in two steps, i.e. in effect applying “local SAM” of Rosseel and Loh (2024) to $\mathbf{Y}^{*}$ given $\mathbf{X}$. Another example of different possibilities in this spirit is provided by Levy (2023) and Levy and McNeish (2025), who carry out both steps using Bayesian estimation.
Another limited-information method that has been developed in the SEM literature is model-implied instrumental variable (MIIV) estimation. It was originally proposed for standard SEMs (see Bollen 2019 for a summary) but it can also be used with binary and ordinal indicators after applying the $\mathbf{Y}$ to $\mathbf{Y}^{*}$ device explained above (Jin et al. 2021; Fisher and Bollen 2020). MIIV is a promising approach, with the particular strength that the estimates are unsually robust to various misspecifications of the form of the measurement model. What is interesting about it for our focus is that while its procedure is ostensibly very different, MIIV has some of the key conceptual characteristics of two-step estimation. In particular, it too separates measurement and structural models, in that the same measurement model (which in MIIV can be left somewhat implicit) can be combined with different structural models without estimating them together. This connection with our kind of likelihood-based two-step estimation would be an interesting topic for future research.